Suppose the body in Sample 20.0 is projected at 24.5 m/s, 30° above the horizontal. Find the a. time of flight, b. range, and c. Maximum height.
1
Expert's answer
2020-11-27T07:15:48-0500

Let's write the equations of motion of the body in horizontal and vertical directions:


"x=v_0tcos\theta, (1)""y=v_0tsin\theta-\dfrac{1}{2}gt^2, (2)"


here, "x" is the horizontal displacement of the body (or the range), "v_0=24.5\ \dfrac{m}{s}" is the initial velocity of the body, "t" is the time of flight of the body, "\theta=30^{\circ}" is the launch angle, "y" is the vertical displacement of the body (or the height) and "g=9.8\ \dfrac{m}{s^2}" is the acceleration due to gravity.

a) Let's first find the time that the body takes to reach the maximum height from the kinematic equation:


"v_y=v_0sin\theta-gt_{rise},""0=v_0sin\theta-gt_{rise},""t_{rise}=\dfrac{v_0sin\theta}{g}."


Then, we can find the time of flight of the body:


"t=2t_{rise}=\dfrac{2v_0sin\theta}{g},""t=\dfrac{2\cdot 24.5\ \dfrac{m}{s}\cdot sin30^{\circ}}{9.8\ \dfrac{m}{s^2}}=2.5\ s."

b) We can substitute the time of flight of the body into the first equation and find the range of the body:


"x=24.5\ \dfrac{m}{s}\cdot 2.5\ s\cdot cos30^{\circ}=53\ m."

c) Finally, we can substitute "t_{rise}" into the second equation and find the maximum height:


"y_{max}=v_0sin\theta\cdot\dfrac{v_0sin\theta}{g}-\dfrac{1}{2}g(\dfrac{v_0sin\theta}{g})^2,""y_{max}=\dfrac{v_0^2sin^2\theta}{2g},""y_{max}=\dfrac{(24.5\ \dfrac{m}{s})^2\cdot sin^230^{\circ}}{2\cdot 9.8\ \dfrac{m}{s^2}}=7.66\ m."

Answer:

a) "t=2.5\ s."

b) "x=53\ m."

c) "y_{max}=7.66\ m."